OpenStudy (taz007):

Help, please! Four-fifths of my current age is greater than three-quarters of my age one year from now. Four-fifths of my current age is also greater than five-sixths of my age one year ago. Given that my age is an integer, what are all possible values for my age?

OpenStudy (alekos):

have you attempted this?

OpenStudy (campbell_st):

let you current age be x then one year from now is x + 1 so the information shows \[\frac{4}{5} \times x > \frac{3}{4} \times (x + 1)\] hope it helps

OpenStudy (taz007):

let y = my age 4/5y > 3/4y +1 -3/4y + 4/5y > 1 (4.5) -3/4y + 4/5y (4.5) > 1 * (4*5) -15y + 16 y > 20 y > 20

OpenStudy (taz007):

For the second portion: let y = my age 4/5y > 5/6y -1 -5/6y + 4/5y > -1 (6*5) -5/6y + 4/5y (6*5) > -1 (6*5) -25y-24y > -30 -1y > -30 y < 30

OpenStudy (taz007):

So, all of the possible values for my age are 21,22,23,24,25,26,27,28,and 29.

OpenStudy (campbell_st):

that looks correct... your age is between 21 and 30... so \[21 \le age \le 30\]

OpenStudy (taz007):

So it should read 20 < y < 30 (greater than but not equeal to 20 and less than but not equal to 30)?

OpenStudy (anonymous):

Refer to the Mathematica attachment.

OpenStudy (taz007):

Thank you!

OpenStudy (anonymous):

Your welcome.

OpenStudy (alekos):

thats what i got.....between 15 and 25